# Farsight: Fostering Responsible AI Awareness During AI Application Prototyping

### v1

@@ E.1. Example Harms Collected from Participants
Teachers use it to help students identify toxic language. Students may feel frustrated or anxious about the accuracy of the AI tool.
2, 1, 3
2, 1, 4
Victims
@@
Scammers use it to identify potential victims. Victims of scams may lose money or personal information.
2, 2, 1
4, 3, 1
Article Sumamrizer (Harm Envisioning Activities)
Anyone reading the document
@@
If the document is sent to anyone else, the summary may not well-represent the document. The readers of the summary might mis-understand the original document.
4, 3, 3
3, 2, 3
Business
@@
Write marketing content, memo. Bad summary can lead to miscommunication, causing financial loss.
4, 3, 3
3, 3, 4
Company publishing the content (news paper)
@@
If the summary is wrong, people would lose confidence in the company.
3, 4, 4
2, 4, 3
Economic loss
@@
The summary may contain misinformation about some critical article, causing economic loss.
3, 3, 1
3, 3, 1
Employees
@@
Employees can use it to summarize documents in a workplace. Wrong summary can harm the stakeholders.
3, 3, 3
4, 3, 3
End user
Change the meaning of the article.
3, 4, 4
3, 4, 4
Journalists / writers
If the summary is wrong, journalists’ reputation might be harmed.
1, 4, 3
1, 3, 4
Kids
Kids use it cheat in school for assignments.
3, 2, 3
2, 2, 3
Readers
@@
AI fails to pick up key information (nuances) in the article =¿ the readers miss the points from the article, miscommunication.
3, 3, 3
2, 2, 2
Readers, content providers, new
@@
If there are two facts in the article that are true independently, then it’s possible that the summary combines them where the statement is no longer true.
3, 2, 3
2, 2, 3
Readers, society
@@
A lot of content is missed, making readers less educated overtime, causing societal loss (some events being lost in history).
3, 2, 1
3, 3, 1
Readers, writers
It would lose a lot of context and details.
3, 3, 3
2, 3, 2
Social groups
@@
Representational harm. It could stereotype some social groups in the output.
3, 3, 3
3, 3, 3
Students
@@
Students can use this feature to summarize articles instead of reading for assignments.
4, 3, 3
3, 3, 3
Students
Students can use this tool to cheat and lose opportunities to learn.
4, 1, 3
3, 1, 4
Students
@@
Students misuse this feature to cheat on school assignments. They might lose learning opportunities.
4, 2, 3
3, 2, 3
Users
The context can be manipulated.
2, 4, 3
2, 4, 3
Users
@@
The summary can lose some information and context from the article.
3, 4, 3
3, 4, 2
Users
@@
The tool makes mistakes (e.g., wrong summary). It can harm the users of the tool by missing information in the article.
3, 3, 3
3, 3, 3
Users
@@
Users may lose trust if the summary is not right. Users would become more stressful as well.
3, 3, 4
2, 2, 3
Writers
@@
If someone uses the tool to write, it may change people’s perception about how you write.
3, 1, 1
2, 1, 2
Math tutor (Harm Envisioning Activities)
Company of the AI product
@@
Engineers use it to explain to non-technical stakeholders. The company may face increased legal liability due to AI-generated explanations being inaccurate or misleading.
2, 1, 2
1, 1, 2
Editors
@@
Journal readers use it to get explanations of equations in inferential statistics sections. Journal editors may worry that the quality of their journal declines because the AI feature makes too many errors. Readers get frustrated over time.
2, 1, 2
2, 1, 3
Marginalized population
@@
People with less resources are less likely to access this tool, losing opportunity to learn.
4, 3, 3
3, 3, 3
Minority social groups
@@
If you ask the math tool about important social groups, it can refuse to answer =¿ ignoring important questions, marginalizing some social groups.
3, 3, 1
3, 3, 1
Minority social groups
@@
There are different ways to phrase things differently based on some social groups. If the user asks a question in non-profession English or non-English, it can cause alienation.
4, 3, 3
3, 3, 3
Non-technical stakeholder
@@
Engineers use it to explain complex math concepts to non-technical stakeholders. Non-technical stakeholders may be misled by AI-generated explanations of complex mathematical models.
3, 2, 3
3, 2, 3
Parents of students
@@
Students use it to learn math. Parents may feel frustrated or anxious about their child’s education.
2, 2, 3
2, 2, 3
Public
The public may be misled by AI-generated explanations of scientific concepts.
3, 3, 3
3, 3, 3
Rocket engineers
Wrong answers can harm the task.
3, 3, 3
3, 2, 3
Student
@@
Tutors use it to help students understand math concepts. Students may lose the opportunity to learn the material in a way that is tailored to their individual needs.
4, 2, 4
4, 2, 4
Students
@@
If the students use this tool to cheat, they would have economic loss (not performing well on their jobs).
2, 1, 3
2, 1, 3
Students
@@
Incorrect math answers harm the user’s learning process.
3, 3, 3
2, 3, 3
Students
Students use it to cheat, losing opportunities to learn.
4, 1, 3
3, 1, 3
Students
@@
Students use it to learn math. Students may feel like they are losing their ability to learn about math problems.
2, 1, 2
2, 1, 2
Students
@@
Students use it to learn math. Students who do not use this AI product may feel like they are not getting the same quality of education as their peers.
3, 3, 3
2, 3, 3
Students
@@
Wrong solution, less optimized solution, wrong information, students learn the wrong thing.
3, 3, 3
3, 2, 3
Students and teachers
Students can use this app to cheat. Students do not learn.
4, 3, 3
3, 2, 4
Students who do not use this AI product
@@
Students use it to learn math. Students who do not use this AI product may feel left behind by their peers who do.
3, 3, 4
3, 2, 4
Students, society
@@
If the answer is based on some context of the math problem, it can harm minority students.
2, 3, 1
3, 3, 1
Users
@@
If you ask tax rates / vote count / interest rate / wages, the wrong answer can cause political harm and social harm.
3, 2, 3
3, 2, 3
Users
@@
Not all questions are about numbers. Sometimes the app might refuse to answer hard math questions, and uses would feel upset.
3, 3, 3
2, 2, 3

### v2

@@ E.1. Example Harms Collected from Participants
Social media users Social media users may feel frustrated or anxious about the accuracy of the AI tool. 3, 3, 3 2, 4, 2
Social media users Social media users who do not use this AI might be exposed to some distorted message. 2, 1, 3 3, 1, 3
Social media users Unfairly flagged as toxic (false positive trigger). 3, 3, 3 3, 2, 2
Student Teachers use it to help students identify toxic language. Students may feel frustrated or anxious about the accuracy of the AI tool. 2, 1, 3 2, 1, 4
@@
Victims Scammers use it to identify potential victims. Victims of scams may lose money or personal information. 2, 2, 1 4, 3, 1
@@
Article Sumamrizer (Harm Envisioning Activities) Anyone reading the document If the document is sent to anyone else, the summary may not well-represent the document. The readers of the summary might mis-understand the original document. 4, 3, 3 3, 2, 3
@@
Business Write marketing content, memo. Bad summary can lead to miscommunication, causing financial loss. 4, 3, 3 3, 3, 4
@@
Company publishing the content (news paper) If the summary is wrong, people would lose confidence in the company. 3, 4, 4 2, 4, 3
@@
Economic loss The summary may contain misinformation about some critical article, causing economic loss. 3, 3, 1 3, 3, 1
@@
Employees Employees can use it to summarize documents in a workplace. Wrong summary can harm the stakeholders. 3, 3, 3 4, 3, 3
End user Change the meaning of the article. 3, 4, 4 3, 4, 4
Journalists / writers If the summary is wrong, journalists’ reputation might be harmed. 1, 4, 3 1, 3, 4
Kids Kids use it cheat in school for assignments. 3, 2, 3 2, 2, 3
@@
Readers AI fails to pick up key information (nuances) in the article =¿ the readers miss the points from the article, miscommunication. 3, 3, 3 2, 2, 2
@@
Readers, content providers, new If there are two facts in the article that are true independently, then it’s possible that the summary combines them where the statement is no longer true. 3, 2, 3 2, 2, 3
@@
Readers, society A lot of content is missed, making readers less educated overtime, causing societal loss (some events being lost in history). 3, 2, 1 3, 3, 1
Readers, writers It would lose a lot of context and details. 3, 3, 3 2, 3, 2
@@
Social groups Representational harm. It could stereotype some social groups in the output. 3, 3, 3 3, 3, 3
@@
Students Students can use this feature to summarize articles instead of reading for assignments. 4, 3, 3 3, 3, 3
Students Students can use this tool to cheat and lose opportunities to learn. 4, 1, 3 3, 1, 4
@@
Students Students misuse this feature to cheat on school assignments. They might lose learning opportunities. 4, 2, 3 3, 2, 3
Users The context can be manipulated. 2, 4, 3 2, 4, 3
@@
Users The summary can lose some information and context from the article. 3, 4, 3 3, 4, 2
@@
Users The tool makes mistakes (e.g., wrong summary). It can harm the users of the tool by missing information in the article. 3, 3, 3 3, 3, 3
@@
Users Users may lose trust if the summary is not right. Users would become more stressful as well. 3, 3, 4 2, 2, 3
@@
Writers If someone uses the tool to write, it may change people’s perception about how you write. 3, 1, 1 2, 1, 2
@@
Math tutor (Harm Envisioning Activities) Company of the AI product Engineers use it to explain to non-technical stakeholders. The company may face increased legal liability due to AI-generated explanations being inaccurate or misleading. 2, 1, 2 1, 1, 2
@@
Editors Journal readers use it to get explanations of equations in inferential statistics sections. Journal editors may worry that the quality of their journal declines because the AI feature makes too many errors. Readers get frustrated over time. 2, 1, 2 2, 1, 3
@@
Marginalized population People with less resources are less likely to access this tool, losing opportunity to learn. 4, 3, 3 3, 3, 3
@@
Minority social groups If you ask the math tool about important social groups, it can refuse to answer =¿ ignoring important questions, marginalizing some social groups. 3, 3, 1 3, 3, 1
@@
Minority social groups There are different ways to phrase things differently based on some social groups. If the user asks a question in non-profession English or non-English, it can cause alienation. 4, 3, 3 3, 3, 3
@@
Non-technical stakeholder Engineers use it to explain complex math concepts to non-technical stakeholders. Non-technical stakeholders may be misled by AI-generated explanations of complex mathematical models. 3, 2, 3 3, 2, 3
@@
Parents of students Students use it to learn math. Parents may feel frustrated or anxious about their child’s education. 2, 2, 3 2, 2, 3
Public The public may be misled by AI-generated explanations of scientific concepts. 3, 3, 3 3, 3, 3
Rocket engineers Wrong answers can harm the task. 3, 3, 3 3, 2, 3
@@
Student Tutors use it to help students understand math concepts. Students may lose the opportunity to learn the material in a way that is tailored to their individual needs. 4, 2, 4 4, 2, 4
@@
Students If the students use this tool to cheat, they would have economic loss (not performing well on their jobs). 2, 1, 3 2, 1, 3
@@
Students Incorrect math answers harm the user’s learning process. 3, 3, 3 2, 3, 3
Students Students use it to cheat, losing opportunities to learn. 4, 1, 3 3, 1, 3
@@
Students Students use it to learn math. Students may feel like they are losing their ability to learn about math problems. 2, 1, 2 2, 1, 2
@@
Students Students use it to learn math. Students who do not use this AI product may feel like they are not getting the same quality of education as their peers. 3, 3, 3 2, 3, 3
@@
Students Wrong solution, less optimized solution, wrong information, students learn the wrong thing. 3, 3, 3 3, 2, 3
Students and teachers Students can use this app to cheat. Students do not learn. 4, 3, 3 3, 2, 4
@@
Students who do not use this AI product Students use it to learn math. Students who do not use this AI product may feel left behind by their peers who do. 3, 3, 4 3, 2, 4
@@
Students, society If the answer is based on some context of the math problem, it can harm minority students. 2, 3, 1 3, 3, 1
@@
Users If you ask tax rates / vote count / interest rate / wages, the wrong answer can cause political harm and social harm. 3, 2, 3 3, 2, 3
@@
Users Not all questions are about numbers. Sometimes the app might refuse to answer hard math questions, and uses would feel upset. 3, 3, 3 2, 2, 3